Sol Summary – February 21st

Crew 189 Sol Summary Report 21 Feb 2018

Sol 3

Summary Title: First results

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
Normal start: everybody is now accustomed to the physical training even if some legs are already heavy. The third EVA was supposed to deploy the LOAC but it didn’t go that well. The power system had an issue and wires melted once I pluged it in. We then brought the equipement back to the hab, and went on for the EVA, going East to do a bit of exploration near Cactus Road, and shoot photos for the school.
This afternoon was dense: after lunch break, we answered to a French journalist interview, played the first session of Gabriel’s human factor experiment, installed a new shade on the GreenHab, worked on experiments, so that we didn’t have enough time to fix the backpack’s fans. Tomorrow EVA will then be done using the new version, before we take time to take care of these (this time, I really want to do it tomorrow).

Look Ahead Plan: Third deployment try of LOAC, data gathering of MegaARES, exploration of Candor Chasma, backpack’s fans inspection, end of shade replacement in greenhab.

Anomalies in work: Broken light in the bathroom (no replacement one found). Water leak in the Hab Lower deck, broken fans in spacesuits.

Weather: Cold and Cloudy

Crew Physical Status: Fine, already a bit tired, but everybody naped.

EVA: Second

Reports to be filed:
Greenhab Officer

Support Requested: New light for bathroom (see Engineering report)

Copyright © The Mars Society. All rights reserved. | Main Site